In travelling the 70 cm along a rifle barrel, a bullet uniformly accelerates from rest to a velocity of 210 m s-1. Find the acceleration involved and the time taken for which the bullet is in the barrel.


1
Expert's answer
2022-03-20T18:50:13-0400

The acceleration of a bullet

"a=\frac{v_f^2-v_i^2}{2d}=\frac{210^2-0^2}{2\times 0.70}=315\:\rm m/s^2"

The time of motion through the barrel

"t=\frac{v_f-v_i}{a}=\frac{210-0}{315}=0.67\:\rm s"